# Ensemble inequivalence and phase transitions in unlabeled networks

### v1

@@ Abstract
We discover a first-order phase transition in the canonical ensemble of random unlabeled networks with a prescribed average number of links. The transition is caused by the nonconcavity of microcanonical entropy. Above the critical point, the canonical and microcanonical ensembles are equivalent and have a well-behaved thermodynamic limit. Below the critical point, the ensemble equivalence is broken, and the canonical ensemble is a mixture of phases: empty networks and networks with logarithmic average degrees. As a consequence, networks with bounded average degrees do not survive in the thermodynamic limit, decaying into the empty phase. The celebrated percolation transition in labeled networks is thus absent in unlabeled networks. In view of these differences between labeled and unlabeled ensembles, the question of which one should be used as a null model of different real-world networks cannot be ignored.
The Erdős-Rényi (ER) model [1] is truly the “harmonic oscillator” of network science and random graphs. It appears as the first prominent subject in many recognizable textbooks in these areas [5, 6, 2, 3, 4], which is not surprising since it has been used as a Petri dish to study a broad spectrum of phenomena in statistical physics, network science, social science, computer science, graph theory, probability, statistics, and many other disciplines. In the microcanonical ER ensemble, $m$ links connect random pairs of $n$ distinct labeled nodes, while in the canonical ensemble, every pair of those nodes is linked independently with probability $p$. The result in both cases is the maximum-entropy distribution over the space of labeled networks with $n$ nodes and a given exact number of links $m$, or expected number of links $\left\langle{m}\right\rangle=pn(n-1)/2$ [7]. Of particular interest to statistical physics are the ensemble equivalence and critical phenomena, which have been explored at great depths in the ER model [11, 13, 8, 9, 10, 12]. A succinct summary is that the percolation transition is at the average degree ${k}_{p}\equiv 2m/n=1$, the network connectivity and symmetry transition is at ${k}_{c}=\log n$, and the canonical and microcanonical ER ensembles are equivalent for any $p\in[0,1]$ and $m=\left\langle{m}\right\rangle$, according to all definitions of ensemble equivalence [14].
@@
Here, we show that the unlabeled Erdős-Rényi model is entirely different from its well-studied labeled counterpart in the sparse regime with average degree ${k}\ll\log n$. The main cause of this difference is the nonconcavity of entropy, a well-studied origin of ensemble inequivalence [14] and phase transitions [27, 26]. In contrast with the standard labeled ER model, the microcanonical unlabeled ER entropy is not concave; it has an inflection point at ${k}_{c}\approx\log n$. We show that this nonconcavity causes a first-order phase transition at $k=k_{c}$. In the dense regime ${k}\gg{k}_{c}$, the microcanonical and canonical unlabeled ER models are equivalent, while away from that region, this ensemble equivalence is broken. The canonical networks with expected average degree ${k}\ll{k}_{c}$ are just empty with high probability $P\approx 1-{{k}}/2{k}_{c}$, while with probability $1-P$, they have average degree $\approx 2{k}_{c}$. This degeneracy is related to but different from the one in microcanonical networks, which are known to exhibit an extreme form of phase separation at ${k}\ll{k}_{c}$ [15]: most nodes are isolated (have degree 00), while all the links condense within an infinitesimal connected component whose size and average degree are $\approx n{k}/{k}_{c}$ and $\approx{k}_{c}$, respectively. Since $P\to 1$ if ${k}\ll{k}_{c}$, unlabeled canonical ER networks with expected average degrees between 00 and ${k}_{c}$ are in fact not realizable in the thermodynamic limit. A remarkable consequence of this is that the celebrated percolation transition at ${k}_{p}=1$ is entirely absent in the unlabeled Erdős-Rényi model.
Microcanonical entropy.— To obtain these results, we first show that the unlabeled microcanonical ER entropy is not concave. This entropy is $S_{nm}=\log g_{nm}$, where $g_{nm}$ is the number of unlabeled graphs with $n$ nodes and $m$ links [28]. In the case of labeled graphs, the corresponding number is simply (${N\choose m}$ with $N\equiv n(n-1)/2$, and the entropy is everywhere concave. The combinatorics of unlabeled graphs is considerably more intricate, since one unlabeled graph corresponds to a class of isomorphic labeled graphs, i.e., all labeled graphs related to each other by vertex renumbering [29]. The standard solution proceeds via Pólya counting theory [30], and we provide a sketch of the derivation in the Appendix. We show there that the generating function for the number of unlabeled graphs, defined as
@@
where $[r,s]$ and $(r,s)$ are the least common multiple and the greatest common divisor of $r$ and $s$, while $\sum_{\{j_{k}\}}$ stands for summation over all sequences of nonnegative integers $\{j_{k}\}$ such that $\sum_{k=1}^{n}kj_{k}=n$ [31]. We then employ representation (2) to compute $g_{nm}$ numerically, and use these exact numerical results in all the figures in this paper [32]. Yet for analytical purposes, we cannot work with the exact expression (2) directly as it is forbiddingly complicated. Instead, we will rely upon its asymptotics that we discuss next.
Entropy asymptotics and nonconcavity.— The asymptotics for $g_{nm}$ have been worked out in [33], and we summarize the relevant ones below. Notation-wise, if the symbol $\star$’ in $a_{n}\star b_{n}$ is $\ll$’, $\sim$’, $\approx$’, or $\gg$’, it means that $\lim_{n\to\infty}a_{n}/b_{n}=c\geq 0$, and that $c=0$, $0<c<\infty$, $c=1$, and $c=\infty$, respectively.
@@
Observe that (4) implies that, with ${k}-\log n\gg 1$, the number of unlabeled graphs is the number of labeled graphs divided by $n!$. This is consistent with the known fact that labeled ER graphs are fully asymmetric if ${k}\gg\log n$, meaning that any permutation of nodes leads to a different labeled graph with high probability [9, 12], identifying $k_{c}=\log n$ as the symmetry transition point. There exist explicit asymptotics for $g_{nm}$ in the sparse regime ${k}\leq\log n$ as well [33], but we will not need that. Instead, we will rely on (6) in what follows.
@@
We thus see right away that the ensemble equivalence must be broken, if the canonical ensemble is asked to produce graphs with average degree $\left\langle{k}\right\rangle<\log n$. Indeed, since $p_{nm}$ does not have maxima at ${k}<\log n$ other than at ${k}=0$, the contributions of probability mass yielding $\left\langle{k}\right\rangle$ cannot come from a single maximum of $p_{nm}$ in that region. They can come from two maxima, one at ${k}=0$ and another one at some ${k}_{*}>\log n$, and we show below that this is indeed what happens. But then, since the $p_{nm}$ distribution is not dominated by a single maximum at $\left\langle{k}\right\rangle<\log n$, the canonical ensemble with expected average degree $\left\langle{k}\right\rangle$ cannot be approximated by the microcanonical ensemble with the average degree fixed exactly at $\left\langle{k}\right\rangle$.
@@
resulting in significant contributions from $|x|\sim\sqrt{m_{*}}$. Using this Gaussian approximation, we estimate the contribution of the region around $m=m_{*}$ to $Z_{n}$ as $\sqrt{2\pi m_{*}}g_{nm_{*}}q^{m_{*}}=e^{m_{*}}/n!$, the latter equation following from (12). The full estimate of $Z_{n}$, summing the contributions from the $m=0$ and $m=m_{*}$ maxima, is thus
@@
or, with probability $1-P$, they have a number of edges close to $m_{*}$. None of these graphs have the number of edges close to $\langle m\rangle<m_{*}$, see Fig. 2(a). The equivalence with the microcanonical ensemble is thus broken.
@@
In labeled graphs, statistical mixtures of sparser and denser phases are rather common [38, 39, 40] and have been compared to gas-liquid phase transitions [41]. They occur, however, in more sophisticated ensembles that control, in addition to the number of edges, the number of triangles or other small subgraphs [42]. It is remarkable that in unlabeled graphs, condensation phenomena of this kind are seen even in the simplest possible ensemble of random graphs—the Erdős-Rényi graphs—formulated in terms of the number of edges alone.
@@
The fact that the phase transition in unlabeled graphs is at the same point $k_{c}=\log n$ as the graph symmetry transition in labeled graphs [9, 12] is not coincidental. Compared to a maximum-entropy distribution over labeled graphs, the corresponding distribution over unlabeled graphs can be seen as shifting the probability masses from more asymmetric to more symmetric graphs, thus making more symmetric graphs more likely [34]. But then, as soon as the graph density is lowered from high values, where almost all graphs are completely asymmetric [12], to the point $k_{c}=\log n$ that allows graphs to be symmetric, also allowing for the presence nodes of degree 00, what is the simplest way to accommodate the shifted probability masses? As our results indicate, the simplest, maximum-entropy way to do so is just to move these masses to empty graphs, simply because they are maximally symmetric. In the microcanonical ensemble, the corresponding solution is to have a needed fraction of 00-degree nodes—“poor loners” in the social network language—since they contribute hyperexponentially to the amount of graph symmetry. It is remarkable that as soon as the expected average degree is $k\ll k_{c}$, these loners dominate in the thermodynamic limit of both the microcanonical and canonical ensembles.
@@
It would be interesting to see if similar critical phenomena are present in more sophisticated/realistic unlabeled network models, including the stochastic block model, latent space models, and the configuration model. Nothing is known about the unlabeled stochastic block model. First results concerning latent space models were obtained in [34], showing that sparse labeled and unlabeled one-dimensional random geometric graphs are entirely different even in terms of their entropy scaling. The existing results concerning the unlabeled configuration model say that if all nodes have the same degree [43], or even if the degree distribution is broad but the fractions of high- and low-degree nodes are sufficiently small [44, 45], then there are no new interesting critical phenomena: the graphs are fully asymmetric, so the entropy of unlabeled graphs is the entropy of labeled graphs minus $\log n!$. In contrast, a ubiquitous feature of many real-world networks are large fractions of high-degree hubs and low-degree nodes [5, 6], bolstering the network symmetry since star graphs are second to empty graphs in terms of the amount of symmetry. Does this feature of real-world networks lead to critical phenomena akin to those we have established here?
@@ Acknowledgements.
Appendix: Pólya counting theory
@@
Here, we outline the calculations in the spirit of Pólya counting theory that lead to the explicit form of the generating function for counting the unlabeled graphs:

### v2

@@ Abstract
We discover a first-order phase transition in the canonical ensemble of random unlabeled networks with a prescribed average number of links. The transition is caused by the nonconcavity of microcanonical entropy. Above the critical point coinciding with the graph symmetry phase transition, the canonical and microcanonical ensembles are equivalent and have a well-behaved thermodynamic limit. Below the critical point, the ensemble equivalence is broken, and the canonical ensemble is a mixture of phases: empty networks and networks with average degrees diverging logarithmically with the network size. As a consequence, networks with bounded average degrees do not survive in the thermodynamic limit, decaying into the empty phase. The celebrated percolation transition in labeled networks is thus absent in unlabeled networks. In view of these differences between labeled and unlabeled ensembles, the question of which one should be used as a null model of different real-world networks cannot be ignored.
The Erdős-Rényi (ER) model [1] is truly the “harmonic oscillator” of network science and random graphs. It appears as the first prominent subject in many recognizable textbooks in these areas [5, 6, 2, 3, 4], which is not surprising since it has been used as a Petri dish to study a broad spectrum of phenomena in statistical physics, network science, social science, computer science, graph theory, probability, statistics, and many other disciplines. In the microcanonical ER ensemble, $m$ links connect random pairs of $n$ distinct labeled nodes, while in the canonical ensemble, every pair of those nodes is linked independently with probability $p$. The result in both cases is the maximum-entropy distribution over the space of labeled networks with $n$ nodes and a given exact number of links $m$, or expected number of links $\left\langle{m}\right\rangle=pn(n-1)/2$ [7, 8]. Of particular interest to statistical physics are the ensemble equivalence and critical phenomena, which have been explored at great depths in the ER model [9, 10, 11, 12, 13, 14]. A succinct summary is that the percolation transition is at the average degree ${k}_{p}\equiv 2m/n=1$, the network connectivity and symmetry transition [15] is at ${k}_{c}=\log n$, and the canonical and microcanonical ER ensembles are equivalent for any $p\in[0,1]$ and $m=\left\langle{m}\right\rangle$, according to all definitions of ensemble equivalence [16].
@@
Here, we show that the unlabeled Erdős-Rényi model is entirely different from its well-studied labeled counterpart in the sparse regime with average degree ${k}\ll\log n$. The main cause of this difference is the nonconcavity of entropy, a well-studied origin of ensemble inequivalence [16] and phase transitions [32, 31, 33, 34]. In contrast with the standard labeled ER model, the microcanonical unlabeled ER entropy is not concave; it has an inflection point at ${k}_{c}\approx\log n$. We show that this nonconcavity causes a first-order phase transition at $k=k_{c}$. In the dense regime ${k}\gg{k}_{c}$, the microcanonical and canonical unlabeled ER models are equivalent, while away from that region, this ensemble equivalence is broken. The canonical networks with expected average degree ${k}\ll{k}_{c}$ are just empty with high probability $P\approx 1-{{k}}/2{k}_{c}$, while with probability $1-P$, they have average degree $\approx 2{k}_{c}$. This degeneracy is related to but different from the one in microcanonical networks, which are known to exhibit an extreme form of phase separation at ${k}\ll{k}_{c}$ [17]: most nodes are isolated (have degree 00), while all the links condense within an infinitesimal connected component whose size and average degree are $\approx n{k}/{k}_{c}$ and $\approx{k}_{c}$, respectively. Since $P\to 1$ if ${k}\ll{k}_{c}$, unlabeled canonical ER networks with expected average degrees between 00 and ${k}_{c}$ are in fact not realizable in the thermodynamic limit. A remarkable consequence of this is that the celebrated percolation transition at ${k}_{p}=1$ is entirely absent in the unlabeled Erdős-Rényi model.
Microcanonical entropy.— To obtain these results, we first show that the unlabeled microcanonical ER entropy is not concave. This entropy is $S_{nm}=\log g_{nm}$, where $g_{nm}$ is the number of unlabeled graphs with $n$ nodes and $m$ links [35]. In the case of labeled graphs, the corresponding number is simply (${N\choose m}$ with $N\equiv n(n-1)/2$, and the entropy is everywhere concave. The combinatorics of unlabeled graphs is considerably more intricate, since one unlabeled graph corresponds to a class of isomorphic labeled graphs, i.e., all labeled graphs related to each other by vertex renumbering [36]. The standard solution proceeds via Pólya counting theory [37], and we provide a sketch of the derivation in the Supplemental Material [25]. The generating function for the number of unlabeled graphs, defined as
@@
where $\mathrm{LCM}(r,s)$ and $\mathrm{GCD}(r,s)$ are the least common multiple and the greatest common divisor of $r$ and $s$, while $\sum_{\{j_{k}\}}$ stands for summation over all sequences of nonnegative integers $\{j_{k}\}$ such that $\sum_{k=1}^{n}kj_{k}=n$ [38]. We then employ representation (2) to compute $g_{nm}$ numerically, and use these exact numerical results in all the figures in this paper [39]. Yet for analytical purposes, we cannot work with the exact expression (2) directly as it is forbiddingly complicated. Instead, we will rely upon its asymptotics that we discuss next.
Entropy asymptotics and nonconcavity.— The asymptotics for $g_{nm}$ have been worked out in [40], and we summarize the relevant ones below. Notationwise, if the symbol $\star$’ in $a_{n}\star b_{n}$ is $\ll$’, $\sim$’, $\approx$’, or $\gg$’, it means that $\lim_{n\to\infty}a_{n}/b_{n}=c\geq 0$, and that $c=0$, $0<c<\infty$, $c=1$, and $c=\infty$, respectively.
@@
Observe that (4) implies that if ${k}-\log n\gg 1$, then the number of unlabeled graphs is the number of labeled graphs divided by $n!$. This is consistent with the known fact that labeled ER graphs are asymmetric [15] if ${k}\gg\log n$, identifying $k_{c}=\log n$ as the symmetry transition point. There exist explicit asymptotics for $g_{nm}$ in the sparse regime ${k}\leq\log n$ as well [40], but we will not need that. Instead, we will rely on (6) in what follows.
@@
We thus see right away that the ensemble equivalence must be broken if the canonical ensemble is asked to produce graphs with average degree $\left\langle{k}\right\rangle<\log n$. Indeed, since $p_{nm}$ does not have maxima at ${k}<\log n$ other than at ${k}=0$, the contributions of probability mass yielding $\left\langle{k}\right\rangle$ cannot come from a single maximum of $p_{nm}$ in that region. They can come from two maxima, one at ${k}=0$ and another one at some ${k}_{*}>\log n$, and we show below that this is indeed what happens. But then, since the $p_{nm}$ distribution is not dominated by a single maximum at $\left\langle{k}\right\rangle<\log n$, the canonical ensemble with expected average degree $\left\langle{k}\right\rangle$ cannot be approximated by the microcanonical ensemble with the average degree fixed exactly at $\left\langle{k}\right\rangle$.
@@
resulting in significant contributions from $|x|\lesssim\sqrt{m_{*}}$. Using this Gaussian approximation, we estimate the contribution of the region around $m=m_{*}$ to $Z_{n}$ as $\sqrt{2\pi m_{*}}g_{nm_{*}}q^{m_{*}}=e^{m_{*}}/n!$, the latter equation following from (12). The full estimate of $Z_{n}$, summing the contributions from the $m=0$ and $m=m_{*}$ maxima, is thus
@@
or, with probability $1-P$, they have a number of edges close to $m_{*}$. None of these graphs have the number of edges close to $\langle m\rangle<m_{*}$, see Fig. 2(a). The equivalence with the microcanonical ensemble is thus broken [44].
@@
Our result that the unlabeled canonical Erdős-Rényi model in the sparse regime is not equivalent to its microcanonical cousin [17] is sad news. If the two models were equivalent, the known properties of the microcanonical model, including its degree distribution and clustering [17], would characterize the canonical model as well, and would not have to be studied “from scratch.” Unfortunately, this is not the case, so even the degree distribution in the canonical model remains unknown, an interesting open question for future work.
@@
In labeled graphs, statistical mixtures of sparser and denser phases are rather common [46, 47, 33] and have been compared to gas-liquid phase transitions [48]. They occur, however, in more sophisticated ensembles that control, in addition to the number of edges, the number of triangles or other small subgraphs [49, 50]. It is remarkable that, in unlabeled graphs, condensation phenomena of this kind are seen even in the simplest possible ensemble of random graphs—the Erdős-Rényi graphs—formulated in terms of the number of edges alone.
@@
The fact that the phase transition in unlabeled graphs is at the same point $k_{c}=\log n$ as the graph symmetry transition in labeled graphs [13, 14] is not coincidental. Compared to a maximum-entropy distribution over labeled graphs, the corresponding distribution over unlabeled graphs can be seen as shifting the probability masses from more asymmetric to more symmetric graphs, thus making more symmetric graphs more likely [41]. But then, as soon as the graph density is lowered from high values, where almost all graphs are asymmetric [14], to the point $k_{c}=\log n$ that allows graphs to be symmetric, also allowing for the presence of nodes of degree 00, what is the simplest way to accommodate the shifted probability masses? As our results indicate, the simplest, maximum-entropy way to do so is just to move these masses to empty graphs, simply because they are maximally symmetric. In the microcanonical ensemble, the corresponding solution is to have a needed fraction of 00-degree nodes—“poor loners” in the social network language—since they contribute hyperexponentially to the amount of graph symmetry. It is remarkable that as soon as the expected average degree is $k\ll k_{c}$, these loners dominate in the thermodynamic limit of both the microcanonical and canonical ensembles.
@@
It would be interesting to see if similar critical phenomena (related to more symmetric graphs being more likely) are present in more sophisticated/realistic unlabeled network models, including the stochastic block model, latent space models, and the configuration model [51, 52]. Nothing is known about the unlabeled stochastic block model. First results concerning latent space models were obtained in [41], showing that sparse labeled and unlabeled one-dimensional random geometric graphs are entirely different even in terms of their entropy scaling. The existing results concerning the unlabeled configuration model say that if all nodes have the same degree [53], or even if the degree distribution is broad but the fractions of high- and low-degree nodes are sufficiently small [54, 55], then there are no new interesting critical phenomena: the graphs are asymmetric, so the entropy of unlabeled graphs is the entropy of labeled graphs minus $\log n!$. In contrast, a ubiquitous feature of many real-world networks are large fractions of high-degree hubs and low-degree nodes [5, 6], bolstering the network symmetry since star graphs are second to empty graphs in terms of the amount of symmetry. Does this feature of real-world networks lead to critical phenomena akin to those we have established here?
@@ Acknowledgements.
SUPPLEMENTAL MATERIAL
Labeled vs. unlabeled network models and their applications
Our focus in the main text has been on presenting surprising thermodynamic features of the simplest possible canonical ensemble of unlabeled graphs. This problem evidently had to be considered before any other unlabeled graph ensembles. If more sophisticated and more realistic ensembles are studied, it will be necessary to keep in mind the features of the simplest ensemble we have analyzed. It is interesting, however, to contemplate potential future applications.
We have remarked in the main text on possible connections to physical and material networks. A majority of models of such networks discussed in the past literature [24, 19, 22, 21, 23, 18, 20] are actually labeled. This is because most of them, but not all, are “mechanistic,” nonequilibrium models, as opposed to null models, e.g., standard canonical ensembles in equilibrium statistical mechanics. Classic examples of mechanistic models include models of absorption and preferential attachment. They tend to describe growing systems in which new elements join the system one at a time, and hence there is a preferred and physically meaningful labeling of elements by their birth times, or more generally, by some other system-specific variables.
The equilibrium null models in statistical mechanics are of very different nature. They do not attempt to reproduce or describe any physical dynamics of particles in the system. Instead, they specify the probability to find the system in a particular microscopic state, given all the macroscopic information about the system. All of the standard null models in equilibrium statistical mechanics, including the microcanonical, canonical, and grand canonical ensembles, can be derived from the maximum entropy principle [26], which simply postulates that the probability distributions over microstates in the ensembles maximize the Shannon entropy of the distribution subject to the constrains coming from the macroscopic information about the system. This means that these models are actually defined by direct specification of probability distributions over microstates. In network models, these microstates are individual graphs, and the macroscopic information is, for instance, the number of edges in the Erdős-Rényi model.
The labeled Erdős-Rényi model admits both interpretations. Mechanistically, it is a model in which (${n\choose 2}$ distinguishable pairs of $n$ distinguishable elements are connected independently with probability $p$. This can be formulated as a growth model, say, in terms of adding new edges one at a time. Yet, the model is also a standard canonical ensemble defined as the probability distribution (over the space of labeled graphs on $n$ vertices) that maximizes the Shannon entropy, subject to the constraint that the average number of edges in these graphs is (${n\choose 2}p$ [7].
If the microscopic elements are now indistinguishable, like molecules and atoms in physical and material networks, then we believe it is obvious that null models associated with such networks must be models of unlabeled graphs. In case of the Erdős-Rényi model discussed here, its unlabeled version is still defined as an entropy-maximizing distribution, except over unlabeled graphs, but as opposed to the labeled Erdős-Rényi model, its unlabeled version does not immediately suggest a mechanistic interpretation.
We note that using labeled null models to analyze networks with unlabeled nodes would lead to paradoxes similar to the (in)famous Gibbs paradox: treating indistinguishable particles as distinguishable leads to an incorrect, nonextensive expression for the entropy of the ideal gas. See [27, 28] and references therein for the coverage of the dramatic 150-year-long history of the Gibbs paradox, and the role that quantum mechanics does (not!) play there.
In applications to physical networks, the unlabeled Erdős-Rényi model is probably not too realistic. One can imagine, in principle, disordered systems of molecules or atoms in which the expected number of connections—contacts or other interactions—is controlled by some external parameters like pressure. If such a system is described by the Erdős-Rényi model, then in view of the results presented here, the difference between labeled and unlabeled behaviors is drastic and can be easily detected in experiments. Indeed, as the pressure parameter is lowered from above to the critical point corresponding to the average degree $k_{c}=\log n$, nothing horrible happens to the labeled system: such a “gel” consisting of distinguishable particles acquires some small disconnected components, but the giant component is still there. If the particles are indistinguishable, however, then the gel “evaporates”—the corresponding graph is empty with high probability.
We do not have an obvious example of a real-world system of particles for which the Erdős-Rényi model would be an appropriate null model, since particles in such systems exist in space, and spatial constraints play an important role. That is why a natural first step to make the unlabeled Erdős-Rényi model more realistic in application to physical networks is to make it spatial, e.g., by considering random geometric graphs (RGGs), mentioned in [19, 18, 20]. We reiterate, however, that these references deal with labeled RGGs, without mentioning potential differences with unlabeled RGGs, and that it was shown in [41] that even such a basic thermodynamic characteristic as the leading term of entropy is different in labeled versus unlabeled one-dimensional RGGs. It would be interesting to see these or similar differences confirmed in real-world experiments with physical networks, and our considerations of the simple unlabeled Erdős-Rényi ensemble are an essential prerequisite for such future studies.
Pólya counting theory
@@
For the convenience of the readers, we outline here the calculations in the spirit of Pólya counting theory that lead to the explicit form of the generating function for counting the unlabeled graphs: